OpenStudy (anonymous):

A right triangle has an angle that measures 34 and the adjacent side measures 17. What is the length of the hypotenuse to the nearest tenth

OpenStudy (♪chibiterasu):

We are including the adjacent and hypotenuse, so we will be using cos. \[\cos 34 = \frac{ adjacent }{ hypostenuse }\] \[\cos 34 = \frac{17}{?}\] Now move the unknown hypotenuse to the other side: \[(x)(\cos 34) = 17\] \[\cos 34 \approx 0.83\] \[(x)(0.83) = 17\] Now divide the cos 34 on both sides: \[\frac{ (x)(0.83) }{ 0.83 } = \frac{ 17 }{ 0.83}\] \[x \approx 20.48\]

OpenStudy (♪chibiterasu):

@Haliemichelle